How do you find the answer to the equation: cubic root of 12 multiplied by the cubic root of 36?
Answer by stanbon(57347)

How do you find the answer to the equation:
cubic root of 12 multiplied by the cubic root of 36?
---
(12)^(1/3)* (36)^(1/3)
---
= (2^2*3 *2^2*3^2)^(1/3)
---
= (2^4*3^3)^(1/3)
---
= 2*3(3)^(1/3)
= 6*(cube root (3))
===========================
Another way to work it:
(12)^(1/3)* (36)^(1/3)
= (6*2*6^2)^(1/3)
----
= (6^3*2)^(1/3)
----
= 6*(cube root(2))
========================
Cheers,
Stan H.